Pls.answer this. Use continuous division to get the prime factorization of each number.
1. 1432
2. 4239
3. 9125
4. 6012
5. 1029
6. 1260
7. 2322
8. 4986

Answer by tommyt3rd(5050)

1. 1432=2*716=2^2*358=2^3*179
2. 4239=3^3*157
3. 9125=5^3*73
4. 6012=2^2*3^2*167
5. 1029=3*7^3
6. 1260=2^2*3^2*5*7
7. 2322=2*3^3*43
8. 4986=2*3^2*277


:)